OpenStudy (anonymous):

Find the derivative of the function. f(t)= sqrt(t/(t^2+1)) I found... (1-t^2)/(2 sqrt(t/(1+t^2)) (1+t^2)^2) is thatright ?

OpenStudy (anonymous):

\[\frac{ d }{dx }\left( \frac{ u }{v } \right)=\frac{ vu'-uv' }{v ^{2} }\]

OpenStudy (anonymous):

so?